# Further author information: (Send correspondence to Chelsea Adelman)

### v1

@@ Acknowledgements.
We would like to thank Cal Bridge for providing financial support for C.A. graduate student research. M.D.F. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303911.References
- [1] Pope, B., Martinache, F., and Tuthill, P., “Dancing in the Dark: New Brown Dwarf Binaries from Kernel Phase Interferometry,” The Astrophysical Journal 767, 110 (Apr. 2013).
- [2] Kammerer, J., Cooper, R. A., Vandal, T., Thatte, D., Martinache, F., Sivaramakrishnan, A., Chaushev, A., Stolker, T., Lloyd, J. P., Albert, L., Doyon, R., Sallum, S., Perrin, M. D., Pueyo, L., Mérand, A., Gallenne, A., Greenbaum, A., Sanchez-Bermudez, J., Blakely, D., Johnstone, D., Volk, K., Martel, A., Goudfrooij, P., Meyer, M. R., Willott, C. J., Furio, M. D., Dang, L., Radica, M., and Noirot, G., “The Near Infrared Imager and Slitless Spectrograph for JWST. V. Kernel Phase Imaging and Data Analysis,” Publications of the Astronomical Society of the Pacific 135, 014502 (Jan. 2023).
@@
- [3] Adelman, C. and Sallum, S., “Planet search around seven white dwarfs in the Hyades cluster using kernel phase interferometry,” in [Optical and Infrared Interferometry and Imaging IX ], Kammerer, J., Sallum, S., and Sanchez-Bermudez, J., eds., Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series 13095, 130952V (Aug. 2024).
- [4] Pope, B., Tuthill, P., Hinkley, S., Ireland, M. J., Greenbaum, A., Latyshev, A., Monnier, J. D., and Martinache, F., “The Palomar kernel-phase experiment: testing kernel phase interferometry for ground-based astronomical observations,” Monthly Notices of the Royal Astronomical Society 455, 1647–1653 (Jan. 2016).
- [5] Kammerer, J., Ireland, M. J., Martinache, F., and Girard, J. H., “Kernel phase imaging with VLT/NACO: high-contrast detection of new candidate low-mass stellar companions at the diffraction limit,” Monthly Notices of the Royal Astronomical Society 486, 639–654 (June 2019).
- [6] Ireland, M. J. and Kraus, A. L., “Orbital Motion and Multi-Wavelength Monitoring of LkCa15 b,” in [Exploring the Formation and Evolution of Planetary Systems ], Booth, M., Matthews, B. C., and Graham, J. R., eds., IAU Symposium 299, 199–203 (Jan. 2014).
- [7] Martinache, F., “Kernel Phase in Fizeau Interferometry,” The Astrophysical Journal 724, 464–469 (Nov. 2010).
- [8] Sallum, S. and Skemer, A., “Comparing nonredundant masking and filled-aperture kernel phase for exoplanet detection and characterization,” Journal of Astronomical Telescopes, Instruments, and Systems 5, 018001 (Jan. 2019).
- [9] Ireland, M. J., “Phase errors in diffraction-limited imaging: contrast limits for sparse aperture masking,” Monthly Notices of the Royal Astronomical Society 433, 1718–1728 (Aug. 2013).
- [10] Vanderburg, A., Rappaport, S. A., Xu, S., Crossfield, I. J. M., Becker, J. C., Gary, B., Murgas, F., Blouin, S., Kaye, T. G., Palle, E., Melis, C., Morris, B. M., Kreidberg, L., Gorjian, V., Morley, C. V., Mann, A. W., Parviainen, H., Pearce, L. A., Newton, E. R., Carrillo, A., Zuckerman, B., Nelson, L., Zeimann, G., Brown, W. R., Tronsgaard, R., Klein, B., Ricker, G. R., Vanderspek, R. K., Latham, D. W., Seager, S., Winn, J. N., Jenkins, J. M., Adams, F. C., Benneke, B., Berardo, D., Buchhave, L. A., Caldwell, D. A., Christiansen, J. L., Collins, K. A., Colón, K. D., Daylan, T., Doty, J., Doyle, A. E., Dragomir, D., Dressing, C., Dufour, P., Fukui, A., Glidden, A., Guerrero, N. M., Guo, X., Heng, K., Henriksen, A. I., Huang, C. X., Kaltenegger, L., Kane, S. R., Lewis, J. A., Lissauer, J. J., Morales, F., Narita, N., Pepper, J., Rose, M. E., Smith, J. C., Stassun, K. G., and Yu, L., “A giant planet candidate transiting a white dwarf,” Nature 585, 363–367 (Sept. 2020).
- [11] Limbach, M. A., Vanderburg, A., MacDonald, R. J., Stevenson, K. B., Jenkins, S., Blouin, S., Rauscher, E., Bowens-Rubin, R., Gallo, E., Mang, J., Morley, C. V., Sing, D. K., O’Connor, C., Venner, A., and Xu, S., “Thermal Emission and Confirmation of the Frigid White Dwarf Exoplanet WD 1856+534 b,” The Astrophysical Journal Letters 984, L28 (May 2025).
- [12] Limbach, M. A., Vanderburg, A., Venner, A., Blouin, S., Stevenson, K. B., MacDonald, R. J., Jenkins, S., Bowens-Rubin, R., Soares-Furtado, M., Morley, C., Janson, M., Debes, J., Xu, S., Kleisioti, E., Kenworthy, M., Butler, P., Crane, J. D., Osip, D., Shectman, S., and Teske, J., “The MIRI Exoplanets Orbiting White dwarfs (MEOW) Survey: Mid-infrared Excess Reveals a Giant Planet Candidate around a Nearby White Dwarf,” The Astrophysical Journal Letters 973, L11 (Sept. 2024).
- [13] Poulsen, S., Debes, J., Cracraft, M., Mullally, S. E., Reach, W. T., Kilic, M., Mullally, F., Albert, L., Thibault, K., Hermes, J. J., Barclay, T., and Quintana, E. V., “A MIRI Search for Planets and Dust around WD 2149+021,” The Astronomical Journal 167, 257 (June 2024).
- [14] Mullally, S. E., Debes, J., Cracraft, M., Mullally, F., Poulsen, S., Albert, L., Thibault, K., Reach, W. T., Hermes, J. J., Barclay, T., Kilic, M., and Quintana, E. V., “JWST Directly Images Giant Planet Candidates Around Two Metal-Polluted White Dwarf Stars,” arXiv e-prints , arXiv:2401.13153 (Jan. 2024).
@@
- [15] Dupuy, T. J., Liu, M. C., and Leggett, S. K., “Discovery of a Low-luminosity, Tight Substellar Binary at the T/Y Transition,” The Astrophysical Journal 803, 102 (Apr. 2015).
@@
- [16] Liu, M. C., Dupuy, T. J., Bowler, B. P., Leggett, S. K., and Best, W. M. J., “Two Extraordinary Substellar Binaries at the T/Y Transition and the Y-band Fluxes of the Coolest Brown Dwarfs,” The Astrophysical Journal 758, 57 (Oct. 2012).
- [17] De Furio, M., Faherty, J. K., Bardalez Gagliuffi, D. C., Gagné, J., Gonzales, E. C., Kiman, R., Kuchner, M., Marocco, F., Merchan, S. A., Rowland, M., Schneider, A. C., Suárez, G., and Vos, J. M., “Discovery of the Second Y+Y Dwarf Binary System: CWISEP J193518.59-154620.3,” arXiv e-prints , arXiv:2508.17176 (Aug. 2025).
@@
- [18] Marois, C., Lafrenière, D., Doyon, R., Macintosh, B., and Nadeau, D., “Angular Differential Imaging: A Powerful High-Contrast Imaging Technique,” The Astrophysical Journal 641, 556–564 (Apr. 2006).
- [19] Laugier, R., Martinache, F., Cvetojevic, N., Mary, D., Ceau, A., N’Diaye, M., Kammerer, J., Lozi, J., Guyon, O., and Lopez, C., “Angular differential kernel phases,” Astronomy and Astrophysics 636, A21 (Apr. 2020).
- [20] Sallum, S., Ray, S., Kammerer, J., Sivaramakrishnan, A., Cooper, R., Greebaum, A. Z., Thatte, D., De Furio, M., Factor, S. M., Meyer, M. R., Stone, J. M., Carter, A., Biller, B., Hinkley, S., Skemer, A., Suárez, G., Leisenring, J. M., Perrin, M. D., Kraus, A. L., Absil, O., Balmer, W. O., Betti, S. K., Boccaletti, A., Bonavita, M., Bonnefoy, M., Booth, M., Bowler, B. P., Briesemeister, Z. W., Bryan, M. L., Calissendorff, P., Cantalloube, F., Chauvin, G., Chen, C. H., Choquet, E., Christiaens, V., Cugno, G., Currie, T., Danielski, C., Dupuy, T. J., Faherty, J. K., Fitzgerald, M. P., Fortney, J. J., Franson, K., Girard, J. H., Grady, C. A., Gonzales, E. C., Henning, T., Hines, D. C., Hoch, K. K. W., Hood, C. E., Howe, A. R., Janson, M., Kalas, P., Kennedy, G. M., Kenworthy, M. A., Kervella, P., Kitzmann, D., Kuzuhara, M., Lagrange, A.-M., Lagage, P.-O., Lawson, K., Lazzoni, C., Lew, B. W. P., Liu, M. C., Liu, P., Llop-Sayson, J., Lloyd, J. P., Lueber, A., Macintosh, B., Manjavacas, E., Marino, S., Marley, M. S., Marois, C., Martinez, R. A., Matthews, B. C., Matthews, E. C., Mawet, D., Mazoyer, J., McElwain, M. W., Metchev, S., Miles, B. E., Millar-Blanchaer, M. A., Molliere, P., Moran, S. E., Morley, C. V., Mukherjee, S., Palma-Bifani, P., Pantin, E., Patapis, P., Petrus, S., Pueyo, L., Quanz, S. P., Quirrenbach, A., Rebollido, I., Redai, J. A., Ren, B. B., Rickman, E., Samland, M., Sargent, B. A., Schlieder, J. E., Schneider, G., Stapelfeldt, K. R., Sutlieff, B. J., Tamura, M., Tan, X., Theissen, C. A., Uyama, T., Vigan, A., Vasist, M., Vos, J. M., Wagner, K., Wang, J. J., Ward-Duong, K., Whiteford, N., Wolff, S. G., Worthen, K., Wyatt, M. C., Ygouf, M., Zhang, X., Zhang, K., Zhang, Z., Zhou, Y., and Zurlo, A., “The JWST Early Release Science Program for Direct Observations of Exoplanetary Systems. IV. NIRISS Aperture Masking Interferometry Performance and Lessons Learned,” The Astrophysical Journal Letters 963, L2 (Mar. 2024).
- [21] Veras, D. and Gänsicke, B. T., “Detectable close-in planets around white dwarfs through late unpacking,” Monthly Notices of the Royal Astronomical Society 447, 1049–1058 (Feb. 2015).
@@
- [22] Martinache, F., Ceau, A., Laugier, R., Kammerer, J., N’Diaye, M., Mary, D., Cvetojevic, N., and Lopez, C., “Kernel-phase analysis: Aperture modeling prescriptions that minimize calibration errors,” Astronomy and Astrophysics 636, A72 (Apr. 2020).

### v2

@@ Acknowledgements.
We would like to thank Cal Bridge for providing financial support for C.A. graduate student research. M.D.F. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303911. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs 4403, 2124, and 1911. All of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via DOI 10.17909/zcr4-9k86. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts.References
@@
- [1] (2024-08) Planet search around seven white dwarfs in the Hyades cluster using kernel phase interferometry. In Optical and Infrared Interferometry and Imaging IX, J. Kammerer, S. Sallum, and J. Sanchez-Bermudez (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 13095, pp. 130952V. External Links: Document Cited by: §1.
- [2] (2025-08) Discovery of the Second Y+Y Dwarf Binary System: CWISEP J193518.59-154620.3. arXiv e-prints, pp. arXiv:2508.17176. External Links: Document, 2508.17176 Cited by: §4, Table 1, §6, §7.
@@
- [3] (2015-04) Discovery of a Low-luminosity, Tight Substellar Binary at the T/Y Transition. The Astrophysical Journal 803 (2), pp. 102. External Links: Document, 1502.04707 Cited by: §1.
- [4] (2013-08) Phase errors in diffraction-limited imaging: contrast limits for sparse aperture masking. Monthly Notices of the Royal Astronomical Society 433 (2), pp. 1718–1728. External Links: Document, 1301.6205 Cited by: §1.
- [5] (2014-01) Orbital Motion and Multi-Wavelength Monitoring of LkCa15 b. In Exploring the Formation and Evolution of Planetary Systems, M. Booth, B. C. Matthews, and J. R. Graham (Eds.), IAU Symposium, Vol. 299, pp. 199–203. External Links: Document Cited by: §1.
- [6] (2023-01) The Near Infrared Imager and Slitless Spectrograph for JWST. V. Kernel Phase Imaging and Data Analysis. Publications of the Astronomical Society of the Pacific 135 (1043), pp. 014502. External Links: Document, 2210.17528 Cited by: §1, §3, §7, §7.
- [7] (2019-06) Kernel phase imaging with VLT/NACO: high-contrast detection of new candidate low-mass stellar companions at the diffraction limit. Monthly Notices of the Royal Astronomical Society 486 (1), pp. 639–654. External Links: Document, 1903.11252 Cited by: §1, §5.
- [8] (2020-04) Angular differential kernel phases. Astronomy and Astrophysics 636, pp. A21. External Links: Document, 2003.04020 Cited by: §5.
- [9] (2025-05) Thermal Emission and Confirmation of the Frigid White Dwarf Exoplanet WD 1856+534 b. The Astrophysical Journal Letters 984 (1), pp. L28. External Links: Document, 2504.16982 Cited by: §1, §7.
- [10] (2024-09) The MIRI Exoplanets Orbiting White dwarfs (MEOW) Survey: Mid-infrared Excess Reveals a Giant Planet Candidate around a Nearby White Dwarf. The Astrophysical Journal Letters 973 (1), pp. L11. External Links: Document, 2408.16813 Cited by: §1, §7.
@@
- [11] (2012-10) Two Extraordinary Substellar Binaries at the T/Y Transition and the Y-band Fluxes of the Coolest Brown Dwarfs. The Astrophysical Journal 758 (1), pp. 57. External Links: Document, 1206.4044 Cited by: §1.
@@
- [12] (2006-04) Angular Differential Imaging: A Powerful High-Contrast Imaging Technique. The Astrophysical Journal 641 (1), pp. 556–564. External Links: Document, astro-ph/0512335 Cited by: §5.
@@
- [13] (2020-04) Kernel-phase analysis: Aperture modeling prescriptions that minimize calibration errors. Astronomy and Astrophysics 636, pp. A72. External Links: Document, 2003.02032 Cited by: §7.
- [14] (2010-11) Kernel Phase in Fizeau Interferometry. The Astrophysical Journal 724 (1), pp. 464–469. External Links: Document, 1009.3933 Cited by: §1, §2.
- [15] (2024-01) JWST Directly Images Giant Planet Candidates Around Two Metal-Polluted White Dwarf Stars. arXiv e-prints, pp. arXiv:2401.13153. External Links: Document, 2401.13153 Cited by: §1, Table 1, §6, §7.
- [16] (2013-04) Dancing in the Dark: New Brown Dwarf Binaries from Kernel Phase Interferometry. The Astrophysical Journal 767 (2), pp. 110. External Links: Document, 1302.6682 Cited by: §1.
- [17] (2016-01) The Palomar kernel-phase experiment: testing kernel phase interferometry for ground-based astronomical observations. Monthly Notices of the Royal Astronomical Society 455 (2), pp. 1647–1653. External Links: Document, 1510.06406 Cited by: §1.
- [18] (2024-06) A MIRI Search for Planets and Dust around WD 2149+021. The Astronomical Journal 167 (6), pp. 257. External Links: Document, 2311.14165 Cited by: §1, §6, §7.
- [19] (2024-03) The JWST Early Release Science Program for Direct Observations of Exoplanetary Systems. IV. NIRISS Aperture Masking Interferometry Performance and Lessons Learned. The Astrophysical Journal Letters 963 (1), pp. L2. External Links: Document, 2310.11499 Cited by: §5.
- [20] (2019-01) Comparing nonredundant masking and filled-aperture kernel phase for exoplanet detection and characterization. Journal of Astronomical Telescopes, Instruments, and Systems 5, pp. 018001. External Links: Document, 1901.01266 Cited by: §1, §5, §7.
- [21] (2020-09) A giant planet candidate transiting a white dwarf. Nature 585 (7825), pp. 363–367. External Links: Document, 2009.07282 Cited by: §1, §7.
- [22] (2015-02) Detectable close-in planets around white dwarfs through late unpacking. Monthly Notices of the Royal Astronomical Society 447 (2), pp. 1049–1058. External Links: Document, 1411.6012 Cited by: §7.